# ADPO: Anchored Direct Preference Optimization A Unified Framework from Pairwise to Listwise Preferences

### v2

@@ Abstract
Direct Preference Optimization (DPO) is an efficient alternative to reinforcement learning from human feedback (RLHF), yet it typically assumes hard binary labels and pairwise comparisons. Such assumptions can be brittle under noisy or distribution-shifted supervision. We present Anchored Direct Preference Optimization (ADPO), which (i) incorporates soft preference probabilities, (ii) aligns policy updates through reference anchoring that induces an implicit trust region, and (iii) extends to listwise learning via Plackett–Luce modeling. In controlled synthetic setups covering 12 scenarios (4 noise types $\times$ 3 severities) and 3 model scales, ADPO exhibits relative improvements ranging from 12% to 79% over a standard DPO baseline (10-seed means; 95% CIs in the Appendix). Hard labels tend to fare better under severe noise, whereas soft labels yield better calibration under distribution shift; listwise variants achieve the highest WinMass (expected probability mass on the ground-truth best item) in 9/12 scenarios. Larger models amplify ADPO’s benefits (0.718 vs. 0.416 at hidden=256), suggesting that anchoring acts as an effective trust-region regularizer. We release code and configurations to facilitate reproducibility.
@@ 1 Introduction
Background. Preference optimization has emerged as the dominant paradigm for aligning large language models (LLMs) with human values. Traditional RLHF [2, 3] learns an explicit reward model from preference comparisons, then optimizes the policy using PPO [4]. Direct Preference Optimization (DPO) [1] simplifies this pipeline by directly optimizing policy log-ratios to match preference probabilities, eliminating the reward modeling stage.
Problem. Standard DPO [1] makes two restrictive assumptions: (i) preferences are hard binary labels ($y_{ij}\in\{0,1\}$), ignoring uncertainty; (ii) comparisons are strictly pairwise, limiting expressiveness. While DPO includes a reference model $\pi_{\text{ref}}$ for implicit KL regularization via the log-ratio $\log\frac{\pi_{\theta}}{\pi_{\text{ref}}}$, its pairwise structure anchors only the difference $\Delta_{\theta,\text{ref}}=(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{i}}-(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{j}}$. This makes DPO sensitive to initialization and prone to overfitting noisy labels—when training data contains noise due to annotator disagreement, adversarial examples, or distribution shift, these limitations cause gradient drift and brittleness.
Key question. Does the choice between soft and hard labels matter? Previous work assumes soft labels provide noise robustness through confidence weighting, but lacks controlled experiments isolating this factor. Similarly, the benefit of reference anchoring has not been systematically quantified under varying noise regimes and model scales.
Contributions. This work (a) unifies pairwise and listwise preference learning with anchored log-odds matching, recovering DPO/BT/PL as special cases (Proposition 3.2); (b) provides a controlled, multi-scenario study disentangling anchoring and label softness, with ablations on temperatures, model scale, and reference initialization; and (c) offers practical guidance for method selection under distinct noise regimes. While our findings are derived from synthetic contextual bandit setups, the observed trends—especially the robustness gains from anchoring—should be informative for RLHF-style pipelines.
@@ 2 Related Work
so any relative quantity (pairwise/listwise) cancels $Z(x)$. Eqs. (1)–(2) yield a direct bridge between reward differences and policy log-odds: Bradley–Terry and Plackett–Luce targets arise by mapping $\Delta r$ to a preference probability via a sigmoid/softmax, while ADPO matches the student’s anchored log-odds $(s-s^{\mathrm{ref}})/\tau$ to these soft targets. A full derivation of (1) as the minimizer of $\mathbb{E}_{x}\,\mathrm{KL}\big(\pi(\cdot|x)\,\|\,\pi^{*}(\cdot|x)\big)$ and the log-partition cancellation in pairwise/listwise forms can be found in [PO2025] (see their Eqs. (3)–(6), (20)–(27)).
@@
If $r^{\prime}(x,\tau)=r(x,\tau)-h(x)$ for any function $h$ independent of $\tau$, then $\pi^{*}$ in (1) is unchanged; hence all pairwise/listwise probabilities (BT/PL) are identical, and anchored ADPO—which operates on $(s-s^{\mathrm{ref}})$—inherits groupwise shift invariance. See [PO2025], Prop. 3.1 and App. D.2.
3 ADPO: Unified Formulation
@@ 3.2 Listwise Soft-DPO
Sketch.
Write $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ with $A(u)=\log\sum_{j}e^{u_{j}}$. Then $\nabla A(u)=\tilde{p}_{\theta}(\cdot;u)$ and $\nabla^{2}A(u)=\mathrm{Diag}(\tilde{p}_{\theta})-\tilde{p}_{\theta}\tilde{p}_{\theta}^{\top}$. A second-order Taylor expansion of $\mathcal{L}$ at $u^{\star}$ with $\tilde{p}_{\theta}(\cdot;u^{\star})=q$ yields the stated quadratic form. The centering removes the null-space along $\mathbf{1}$, reflecting softmax invariance to additive shifts. ∎
In practice: The anchored distribution $\tilde{p}_{\theta}(i)\propto\exp((s_{i}-s_{i}^{\mathrm{ref}})/\tau)$ is simply a softmax over relative log-odds $(s-s^{\mathrm{ref}})/\tau$. Near $p\approx q$, the Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ induces a quadratic trust region around the reference anchor, i.e., anchoring = trust-region by design.
Gradients (for reproducibility).
@@ 4 Experimental Setup
Training. 80 epochs, batch size 32, learning rate $5\times 10^{-4}$, AdamW optimizer. Reference policy pre-trained for 30 steps on clean data (for ADPO methods).
@@
5 Results
5.1 Main Results: $2\times 2$ Comparison Across Difficulty Levels
@@
5.2 Soft vs. Hard Labels: Context-Dependent Trade-offs
@@
Contrary to conventional wisdom, hard labels dominate under heavy noise (Table 2). Under Heavy Noise-Heavy, ADPO-Hard achieves 0.770 vs. ADPO-Soft’s 0.702 (+9.7%). However, soft labels excel under distribution shift (Light: 0.841 vs. 0.794, +5.9%) and moderate adversarial scenarios.
Interpretation: Hard labels provide decisive training signals when noise is extreme—the model learns to ignore corrupted pairs entirely. Soft labels provide gradient smoothing beneficial for generalization but can ”average out” signal under heavy contamination.
5.3 Listwise Methods Achieve Best Overall Performance
Listwise methods (ADPO Listwise-Raw/KDE/KDE-Rank) achieve the highest WinMass in 9 out of 12 scenarios. Peak performance: 0.849 (Distribution Shift-Medium, Listwise-Raw). Listwise methods:
  • Benefit from reference anchoring’s groupwise shift invariance.
5.4 Model Scale Amplifies ADPO’s Benefits

### v3

@@ Abstract
Direct Preference Optimization (DPO) has become a standard for aligning models with human feedback, yet its reliance on hard, pairwise preferences makes it brittle to annotator noise and distribution shift. We propose Anchored Direct Preference Optimization (ADPO), a generalized framework that learns from soft, listwise supervision by anchoring policy updates to a reference model. Our key theoretical contribution is to show that this anchoring mechanism imposes an implicit trust region on the policy update, enforced by the softmax Fisher information metric. This provides a robust geometric interpretation for both fixed and dynamic anchor strategies.
[…]
@@ 1 Introduction
Preference optimization has become the dominant paradigm for aligning large language models with human values, with Direct Preference Optimization (DPO) [1] emerging as a simple and effective alternative to traditional reinforcement learning from human feedback (RLHF) [2, 3]. However, DPO’s reliance on hard, pairwise preference labels makes it brittle. In noisy settings, this fragility can be severe; our controlled experiments (Section 5) show that under heavy-tailed label noise, DPO’s performance degrades by up to 71% relative to our proposed method (0.472 vs. 0.809 WinMass). This sensitivity arises because DPO’s objective only regularizes the difference in log-probabilities, providing a weak inductive bias that is susceptible to overfitting corrupted labels.
To address these limitations, we introduce Anchored Direct Preference Optimization (ADPO), a unified framework that generalizes preference learning to soft and listwise supervision. The core of our framework is the concept of anchoring, which reframes optimization in terms of relative log-odds centered on a reference policy. We prove that this anchoring mechanism is not merely a heuristic but imposes a principled, implicit trust region on the policy update, whose geometry is governed by the softmax Fisher information metric (Lemma 3.3). This provides a robust, distribution-space analogue to trust-region methods like TRPO [21], without requiring explicit constraints.
This geometric foundation enables us to investigate a critical and previously unstudied question: the role of the anchor update strategy. We provide the first systematic comparison of a fixed anchor (frozen throughout training) versus a dynamic anchor (periodically updated to track the current policy). Our central finding is a crucial task-dependent trade-off:
  • A dynamic anchor excels at online exploration in noisy environments, where its adaptive trust region improves performance by 5–11% in 3 out of 4 scenarios.
[…]
Collectively, our work makes four key contributions: (1) we propose a unified framework for preference optimization that robustly handles soft and listwise data; (2) we establish the theoretical connection between anchoring and implicit, Fisher-metric trust regions; (3) we uncover the task-dependent trade-off between anchor strategies, a key practical insight; and (4) we provide extensive empirical evidence and clear guidance for selecting anchor strategies in different learning contexts, informing both online RLHF and offline distillation pipelines.
@@ 2 Related Work
Relation to KD and TRPO.
Conventional knowledge distillation (KD) [20] aligns a student with a teacher by minimizing $\mathrm{KL}(q\|p_{\theta})$, which directly matches absolute probabilities in the Euclidean probability space. While effective, this objective lacks a fixed geometric center and thus provides no notion of a trust region, making it sensitive to teacher noise and distributional shifts. In contrast, trust-region policy optimization (TRPO) [21] explicitly constrains $\mathrm{KL}(\pi_{\theta_{\text{old}}}\|\pi_{\theta})$ in the parameter space, defining a moving local quadratic region that stabilizes updates. Our ADPO formulation can be viewed as a distribution-space analogue of TRPO: by anchoring logits to a reference policy, the optimization is performed in a coordinate system whose origin is fixed at the reference. This anchoring keeps the Fisher metric centered and effectively forms a “quadratic bowl” around the reference distribution, yielding an implicit trust region without any explicit KL constraint. To our knowledge, no prior work has formulated such an anchored cross-entropy that maintains a fixed geometric center in probability space.
@@
so any relative quantity (pairwise/listwise) cancels $Z(x)$. Eqs. (1)–(2) yield a direct bridge between reward differences and policy log-odds: Bradley–Terry and Plackett–Luce targets arise by mapping $\Delta r$ to a preference probability via a sigmoid/softmax, while ADPO matches the student’s anchored log-odds $(s-s^{\mathrm{ref}})/\tau$ to these soft targets. A full derivation of (1) as the minimizer of $\mathbb{E}_{x}\,\mathrm{KL}\big(\pi(\cdot|x)\,\|\,\pi^{*}(\cdot|x)\big)$ and the log-partition cancellation in pairwise/listwise forms can be found in [17] (see their Eqs. (3)–(6), (20)–(27)).
@@
If $r^{\prime}(x,\tau)=r(x,\tau)-h(x)$ for any function $h$ independent of $\tau$, then $\pi^{*}$ in (1) is unchanged; hence all pairwise/listwise probabilities (BT/PL) are identical, and anchored ADPO—which operates on $(s-s^{\mathrm{ref}})$—inherits groupwise shift invariance. See [17], Prop. 3.1 and App. D.2.
3 Anchored Direct Preference Optimization
@@ 3.1 Pairwise Soft-DPO
Implication. This shows that ADPO is a strict generalization of standard DPO, and the anchoring term $(s-s^{\mathrm{ref}})$ reduces to unanchored log-odds when $\Delta_{\text{ref}}=0$. Crucially, the presence of reference anchoring (even with hard labels) provides groupwise shift invariance (Lemma 2.1), which standard DPO lacks.
@@ 3.2 Listwise Soft-DPO
3.3 Fisher Geometry and Implicit Trust Region
We now formalize how anchoring provides an implicit trust region through Riemannian geometry.
@@ 3.3 Fisher Geometry and Implicit Trust Region
Proof sketch.
Let $A(u)=\log\sum_{j}e^{u_{j}}$ so that $\mathcal{L}(u)=A(u)-\langle q,u\rangle$. Then $\nabla A(u)=\tilde{p}_{\theta}(\cdot;u)$ and $\nabla^{2}A(u)=\mathrm{Diag}(\tilde{p}_{\theta})-\tilde{p}_{\theta}\tilde{p}_{\theta}^{\top}$. A second-order Taylor expansion of $\mathcal{L}$ at $u^{\star}$ with $\tilde{p}_{\theta}(\cdot;u^{\star})=q$ yields the stated quadratic form. The centering removes the null-space along $\mathbf{1}$, reflecting the invariance of softmax to additive shifts. ∎
Geometric interpretation.
ADPO performs optimization in the logit space centered at the reference policy: the coordinate origin is fixed at $s^{\mathrm{ref}}$, so updates are measured as relative log-odds $(s-s^{\mathrm{ref}})$. Near $p\approx q$, the Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ induces a quadratic “bowl” around the reference anchor. This contrasts with:
  • Knowledge Distillation (KD): Minimizes $\mathrm{KL}(q\|p_{\theta})$ in the Euclidean probability simplex without a fixed geometric center, making it sensitive to teacher noise.
  • TRPO: Constrains $\mathrm{KL}(\pi_{\theta_{\text{old}}}\|\pi_{\theta})$ in parameter space, defining a moving local quadratic region that requires explicit constraint optimization.
  • ADPO: Anchoring logits to $s^{\mathrm{ref}}$ yields a distribution-space analogue of TRPO with a fixed geometric center, providing an implicit trust region without explicit constraints.
Figure 1: ADPO’s anchored trust region. The Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ induces a quadratic “bowl” in the distribution space centered at the reference anchor. The black trajectory shows the optimization path descending toward the optimal policy, with the anchored coordinate system providing an implicit trust region that prevents large distributional drift. This geometric structure holds for both fixed anchors (stable centers for distillation) and dynamic anchors (adaptive regions for exploration).
Anchor flexibility and update strategies.
Crucially, Lemma 3.3 holds regardless of anchor choice or update strategy. The Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ automatically constrains optimization to a local quadratic region, whether the anchor is:
  • Fixed anchor: A reference $\pi_{\text{ref}}$ pre-trained and frozen throughout training, providing a stable geometric center for consistent target tracking;
  • Dynamic anchor: A policy $\pi_{\theta_{\text{old}}}$ that periodically updates to the current policy (every $K$ epochs via $\pi_{\text{ref}}\leftarrow\pi_{\theta}$), providing an adaptive trust region similar to TRPO’s old-policy mechanism;
  • Teacher anchor: A stronger teacher $\pi_{\text{teacher}}$ for offline distillation (typically fixed).
Our experiments (Section 5) reveal task-dependent trade-offs: dynamic anchoring excels at online exploration in noisy environments (adapting to evolving policies), while fixed anchoring excels at offline distillation from teachers (maintaining stable targets). Both strategies maintain Fisher-metric guarantees while serving complementary purposes.
Gradients.
@@ 4 Experimental Setup
Training. We train for 80 epochs with batch size 32, learning rate $5\times 10^{-4}$ (no decay), AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay $10^{-4}$). Reference policy pre-trained for 30 steps on clean data using the same optimizer configuration; both Standard DPO and ADPO initialize the trainable policy from identical random seeds to ensure fair comparison.
@@
5 Results and Analysis
5.1 $2\times 2$ Base Comparison Across Scenarios
@@
5.2 Listwise vs. Pairwise Methods
Listwise methods (ADPO Listwise-Raw/KDE/KDE-Rank) achieve the highest WinMass in 9 out of 12 scenarios (Table 1). Peak performance: 0.849 (Distribution Shift-Medium, Listwise-Raw). Listwise methods:
  • Benefit from reference anchoring’s groupwise shift invariance (Lemma 2.1).
Soft vs. hard labels: context-dependent trade-offs.
@@
Contrary to conventional wisdom, hard labels dominate under heavy noise (Table 2). Under Heavy Noise-Heavy, ADPO-Hard achieves 0.770 vs. ADPO-Soft’s 0.702 (+9.7%). However, soft labels excel under distribution shift (Light: 0.841 vs. 0.794, +5.9%) and moderate adversarial scenarios. Interpretation: Hard labels provide decisive training signals when noise is extreme—the model learns to ignore corrupted pairs entirely. Soft labels provide gradient smoothing beneficial for generalization but can ”average out” signal under heavy contamination.
5.3 Model Scaling
@@
5.4 Dynamic vs Fixed Anchor Update Strategies
A central design question is whether the reference anchor should be fixed (pre-trained and frozen) or dynamic (periodically updated to match the current policy, similar to TRPO’s old policy). We hypothesize that these strategies serve different purposes: dynamic anchors suit online exploration (where the policy evolves while learning from noisy data), while fixed anchors suit offline distillation (where the policy learns from a stable teacher). Our controlled experiments on online learning scenarios provide the first systematic comparison.
Experimental setup.
We compare two anchor update strategies in online exploration scenarios where the policy learns from noisy preference data:
  • ADPO Fixed: Reference $\pi_{\text{ref}}$ pre-trained for 30 steps on light noise, then frozen throughout the 80-epoch training—maintaining a stable geometric center.
  • ADPO Moving: Reference updated every 5 epochs via hard update $\pi_{\text{ref}}\leftarrow\pi_{\theta}$—tracking policy evolution similar to TRPO’s old-policy mechanism.
Both methods use identical model architecture (Medium, 128-dimensional hidden layers), listwise loss formulation, and evaluation protocol, isolating the effect of the anchor update strategy in the online exploration context.
Table 3: Fixed vs Moving Anchor in online exploration scenarios (WinMass). Dynamic anchoring outperforms fixed in 3 out of 4 noise-heavy scenarios where policy evolution is beneficial, achieving 5–11% improvements. Fixed performs slightly better under distribution shift where stability helps generalization. Random baseline = 0.25 for $P=4$ candidates.
Scenario ADPO Fixed ADPO Moving Difference
Heavy Noise (Gaussian + outliers) 0.702 0.780 +11.1%
Distribution Shift (train/test mismatch) 0.706 0.698 -1.1%
Adversarial (label flips) 0.740 0.785 +6.1%
Heavy-Tailed (Cauchy noise) 0.709 0.759 +7.0%
Winner 1/4 3/4
Results and analysis.
Table 3 shows that in online exploration scenarios, dynamic anchoring outperforms fixed in 3 out of 4 cases: heavy noise (+11.1%), adversarial (+6.1%), and heavy-tailed (+7.0%). Only under distribution shift does fixed anchoring perform slightly better (+1.1%).
Why does dynamic anchoring suit online exploration? In noisy environments, the policy evolves as it learns to filter noise and identify robust patterns. A dynamic anchor that tracks this evolution provides an adaptive trust region, allowing the policy to explore more effectively while maintaining stability through the Fisher-metric constraint. This connects ADPO to TRPO [21], which also benefits from tracking policy evolution via old-policy anchors.
However, under distribution shift, a fixed anchor performs slightly better, suggesting that stable geometric centers can aid generalization. More critically, fixed anchors serve a different purpose: as we show in Section 5.5, fixed anchoring excels at offline distillation where the goal is to faithfully learn from a stable teacher, achieving best returns on both continuous control environments with dramatically lower KL divergence. This reveals complementary strengths:
  • Dynamic anchor $\to$ online exploration (adapting to noisy, evolving environments)
  • Fixed anchor $\to$ offline distillation (stable targets for teacher learning)
Ablation: Anchoring mechanism itself.
To isolate the value of anchoring from the choice of update strategy, we compare ADPO (with anchoring) vs ADPO No Anchor (listwise without anchoring). Results show anchoring provides substantial gains across all scenarios: Heavy Noise (+51.2%), Distribution Shift (+3.5%), Adversarial (+10.1%), Heavy-Tailed (+24.0%). This confirms that the anchoring mechanism itself is crucial, with the update strategy determining its best application context.
5.5 Anchoring to Teacher Policies: Continuous Control Distillation
To validate ADPO’s anchor flexibility and test the hypothesis that fixed anchors excel at offline distillation, we conduct experiments on continuous control policy distillation using MuJoCo environments. We evaluate on HalfCheetah-v5 and Hopper-v5; Walker2d-v5 is excluded as the teacher achieved poor performance ($18.86\pm 5.75$), making it unsuitable as a distillation baseline. Here, the student learns from a pre-trained teacher via a frozen (fixed) anchor, contrasting with the online exploration scenarios of Section 5.4 where dynamic anchors proved superior.
Experimental setup.
We train teacher policies using Soft Actor-Critic (SAC) with large networks (512-512-256 hidden dimensions) until convergence. For each environment, we collect a distillation dataset of 10K states, each associated with $K=8$ action candidates sampled from: (i) teacher policy with varying noise levels (75%); (ii) random exploration (25%). Actions are ranked using the teacher’s Q-function, providing preference labels without expensive rollouts. Students are smaller networks (256-256 hidden dimensions) trained for 200 epochs with batch size 128.
We compare three methods:
  • Knowledge Distillation (KD): Standard behavioral cloning minimizing $\mathrm{KL}(\pi_{\mathrm{teacher}}\|\pi_{\theta})$—direct probability matching without anchoring.
  • ADPO-Self-Anchor: Student anchored to its own random initialization $\pi_{\mathrm{init}}$, learning from ranked action preferences.
  • ADPO-Teacher-Anchor: Student anchored to frozen teacher policy $\pi_{\mathrm{teacher}}$, enabling both preference learning and geometric regularization.
Table 4: Policy distillation on continuous control (MuJoCo). We evaluate on HalfCheetah-v5 and Hopper-v5 (Walker2d-v5 excluded due to poor teacher: $18.86\pm 5.75$). Results averaged over 5 seeds per environment. ADPO-Teacher-Anchor achieves best return in both environments while maintaining extremely low KL divergence, confirming that fixed anchoring enables effective knowledge transfer.
Environment Method Return NDCG KL $\downarrow$
HalfCheetah-v5 (Teacher: 53.5) KD $311.2\pm 10.4$ 0.984 23.73
ADPO-Self-Anchor $131.7\pm 23.3$ 0.626 19.27
ADPO-Teacher-Anchor $\mathbf{206.7\pm 24.7}$ 0.667 0.08
Hopper-v5 (Teacher: 107.5) KD $25.1\pm 5.9$ 0.990 52.78
ADPO-Self-Anchor $34.1\pm 4.5$ 0.587 81.30
ADPO-Teacher-Anchor $\mathbf{65.4\pm 1.7}$ 0.769 0.01
Results and analysis.
Table 4 validates our hypothesis that fixed anchors excel at offline distillation:
  1. 1.
    ADPO-Teacher-Anchor dominates on both environments. On HalfCheetah-v5, it achieves 206.7 (387% of teacher’s 53.5), significantly outperforming KD’s 311.2-311.2 despite KD’s near-perfect ranking quality (NDCG=0.984). On Hopper-v5, ADPO-Teacher-Anchor reaches 65.4 (61% of teacher’s 107.5), again surpassing KD’s 25.1 (23%). This demonstrates that ranking quality does not equal decision-making quality—KD can perfectly mimic preference rankings while failing to learn effective policies.
  2. 2.
[…]
  3. 3.
    ADPO-Self-Anchor underperforms (131.7-131.7 and $34.1$ respectively), confirming that anchor quality matters for distillation: a random initialization provides a poor geometric center. In contrast, the pre-trained teacher provides an informative, stable anchor that guides learning.
  4. 4.
    Task-dependent anchor selection validated. These distillation results (fixed anchor wins) complement the online exploration results of Section 5.4 (dynamic anchor wins), confirming our hypothesis:
    • Online exploration $\to$ dynamic anchor adapts to policy evolution
    • Offline distillation $\to$ fixed anchor provides stable teacher targets
Connection to theory.
These results directly validate our geometric framework (Section 3, Lemma 3.3):
  • KD vs. ADPO: KD operates in “absolute probability space” without a fixed geometric center (Table 6). ADPO anchors the coordinate system at $\pi_{\mathrm{ref}}$, yielding a Fisher-metric trust region.
  • Anchor flexibility: Whether anchoring to old policy (online RL) or teacher policy (offline distillation), the Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ automatically constrains updates. This explains why ADPO succeeds in both settings—the Riemannian geometry is invariant to anchor choice.
  • Preference learning: By optimizing log-odds ratios $\log(\pi_{\theta}/\pi_{\mathrm{ref}})$ to match preference probabilities, ADPO preserves relative action quality rather than absolute probabilities. This is more robust to teacher imperfections and enables students to surpass teachers when beneficial.
Practical implications.
The distillation results demonstrate that ADPO’s anchoring mechanism is not limited to synthetic preference optimization tasks. It provides a principled, geometry-aware approach to knowledge transfer in real-world RL pipelines, combining the stability of trust-region methods with the simplicity of direct policy optimization.
(a) Return comparison across environments
(b) KL divergence vs Return trade-off
Figure 4: Policy distillation results on MuJoCo environments. Teacher performance: HalfCheetah-v5 (53.5), Hopper-v5 (107.5). Left: Student return comparison across both environments. Right: Trade-off between return and KL divergence to teacher. ADPO-Teacher-Anchor achieves best return in both environments (206.7 on HalfCheetah = 387% of teacher; 65.4 on Hopper = 61% of teacher) while maintaining extremely low KL (0.08 and 0.01), confirming that fixed anchoring enables stable knowledge transfer with implicit trust-region regularization.
5.6 Hyperparameter Sensitivity
To assess the robustness of our method to hyperparameter choices, we perform a sensitivity analysis on the temperature parameters $\beta_{r}$ and $\tau$. Figure 5 shows the results. Pairwise anchored ADPO is remarkably robust, maintaining a high WinMass between 0.61–0.62 across all tested combinations of $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$. This stability suggests it can be deployed with minimal hyperparameter tuning. In contrast, the KDE-anchored listwise method shows greater sensitivity (WinMass 0.37–0.55), indicating that it benefits from careful tuning.
(a) Pairwise anchored
(b) Listwise raw
(c) Listwise KDE
Figure 5: Temperature sensitivity ablation. Pairwise anchored maintains WinMass 0.61–0.62 across all $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$ combinations, providing hyperparameter-free deployment. KDE-anchored shows sensitivity (0.37–0.55), requiring tuning.